Pearl Corp. is expected to have an EBIT of $3,400,000 next year. Depreciation, the increase in net working capital, and capital spending are expected to be $160,000, $155,000, and $195,000, respectively. All are expected to grow at 18 percent per year for four years. The company currently has $17,500,000 in debt and 1,350,000 shares outstanding. After Year 5, the adjusted cash flow from assets is expected to grow at 2.5 percent indefinitely. The company’s WACC is 9.1 percent and the tax rate is 21 percent.

What is the price per share of the company's stock? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

EBIT(1-Tax) = 3,400,000(1-21%) = $2,686,000

Add: Depreciation (non-cash) = 160,000

Less: Increase in Working Capital = $155,000

Capital Spending = $195,000

Free cash flow = $2,496,000

Value of firm is equal to the present value of all future free cash flows

= 2,496,000/(1.091) + 2,496,000*1.18/(1.091)^2 + 2,496,000*(1.18)^2/(1.091)^3 + 2,496,000*(1.18)^3/(1.091)^4 + 2,496,000*(1.18)^4/(1.091)^5 + 2,496,000*(1.18)^4(1.025)/(1.091)^5 (9.1%-2.5%)

= $62,085,469.75

Price per share = (62,085,469.75-17,500,000)/1,350,000

= $33.03